The force of the wind blowing on a verticle surface varies jointly as the area of the surface and the square of the velocity. If a wind blowing at 50 miles per hour exerts a force of 75 pounds on a surface of 500 ft squared, how much force will a wind of 75 mph place on a surface of 10 ft squared.

10 ft squared means 10 ft^ i think.
500 ft squared means 500 ft^.

Answer by stanbon(75887)

The force of the wind blowing on a verticle surface varies jointly as the area of the surface and the square of the velocity.
---
f = kAv^2
--------------
If a wind blowing at 50 miles per hour exerts a force of 75 pounds on a surface of 500 ft squared,
----
solve for "k":
50 = k*500*75^2
k = 50/(500*75^2)
k = 0.0000178
------------------------
how much force will a wind of 75 mph place on a surface of 10 ft squared.
f = (0.0000178)(10)*75^2
force = 1 lb
====================
Cheers,
Stan H.